# Inference Time Alignment of Diffusion Models with Evolutionary Algorithms

### v1

@@
Inference Time Alignment of Diffusion Models with Evolutionary Algorithms
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Evolutionary Algorithms for Inference-Time Alignment of Diffusion Models
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Evolutionary Algorithms for Black-Box Inference-Time Alignment of Diffusion Models
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
Inference-Time Alignment of Diffusion Models with Evolutionary Algorithms
Purvish Jajal*
Purdue University
West Lafayette, IN 47907
pjajal@purdue.edu
&Nick John Eliopoulos*
Purdue University
West Lafayette, IN 47907
neliopou@purdue.edu
&Benjamin Shiue-Hal Chou
Purdue University
West Lafayette, IN 47907
chou150@purdue.edu
&George K. Thiruvathukal
Loyola University Chicago
Chicago, IL 60660
gkt@cs.luc.edu
&James C. Davis
Purdue University
West Lafayette, IN 47907
davisjam@purdue.edu
&Yung-Hsiang Lu
Purdue University
West Lafayette, IN 47907
yunglu@purdue.edu
@@ Abstract
Diffusion models are state-of-the-art generative models in various domains, yet their samples often fail to satisfy downstream objectives such as safety constraints or domain-specific validity. Existing techniques for alignment require gradients, internal model access, or large computational budgets. We introduce an inference-time alignment framework based on evolutionary algorithms. We treat diffusion models as black-boxes and search their latent space to maximize alignment objectives. Our method enables efficient inference-time alignment for both differentiable and non-differentiable alignment objectives across a range of diffusion models. On the DrawBench and Open Image Preferences benchmark, our EA methods outperform state-of-the-art gradient-based and gradient-free inference-time methods. In terms of memory consumption, we require $55\%$ to $76\%$ lower GPU memory than gradient-based methods. In terms of running-time, we are $72\%$ to $80\%$ faster than gradient-based methods. We achieve higher alignment scores over 50 optimization steps on Open Image Preferences than gradient-based and gradient-free methods.
footnotetext: * Purvish Jajal and Nick John Eliopoulos contributed equally to this work.
@@ 1 Introduction
Figure 1: Samples generated by our method on Stable Diffusion-3 across three different prompts under various alignment objectives (brightness, saturation, file size). Each row shows the progression over optimization steps $i$, with the corresponding metric values displayed in the top-right corner. The noise in each generation is optimized via the CoSyNE algorithm [13] over 14 optimization steps. Arrows ($\uparrow$/$\downarrow$) indicate whether the metric is being maximized or minimized.
@@
Diffusion models [40, 34, 9] have emerged as state-of-the-art generative models for synthesizing high-quality images, video, and audio [3, 16, 18, 25]. However, the outputs of diffusion models often fail to meet downstream objectives, such as adhering to user preferences [53, 51, 37], safety constraints [28], or domain-specific validity [14, 49]. This problem is referred to as alignment—ensuring that generated samples satisfy external objectives beyond the original maximum-likelihood objective [24, 48, 2, 21, 5, 31, 4]. Alignment is an important concern for many practical applications such as image generation [53, 51, 37] and drug discovery [14, 49].
Methods for diffusion model alignment can be categorized into two types: fine-tuning-based alignment and inference-time alignment. Fine-tuning-based alignment typically relies on supervised fine-tuning or reinforcement learning to adjust model parameters [24]. In contrast, inference-time alignment modifies the sampling process during generation to maximize a reward function (i.e., alignment objective) [45]. These methods aim to be computationally efficient and support arbitrary rewards while maintaining sample quality. However, existing approaches have several drawbacks: they are often not black-box, requiring access to gradients or internal model states; they cannot efficiently generate many aligned samples; and they are often not sample-efficient. As a result, they are difficult to integrate with a variety of models, sampling strategies, and reward functions; generating many aligned samples becomes cost-prohibitive; and obtaining aligned samples may take too long.
In this work, we demonstrate that evolutionary algorithms (EAs)—specifically genetic algorithms (GA) and evolutionary strategies (ES)—can be used to align diffusion models. As a class of algorithms, EAs have desirable properties: they are black-box, gradient-free optimizers, that can be highly parallelized [39, 36]. By interpreting the diffusion models as black-boxes and searching over their latent space by using EAs we can optimize the alignment objective. Our approach enables efficient alignment that supports both differentiable and non-differentiable rewards, and is compatible with various diffusion models. Fig. 1 depicts our method applied to various alignment objectives.
To quantify alignment quality, GPU-memory footprint, and runtime efficiency, we evaluate our inference-time alignment method on two prompt collections [35, 6], four reward functions [53, 51, 32, 4], and three diffusion models [34, 9, 3]. On both prompt collections, our method outperforms existing gradient-free and gradient-based inference-time alignment approaches. In terms of computational costs, we exhibit $55\%$ to $76\%$ reduction in GPU memory than gradient-based methods, while only consuming up to 1.5×1.5$\times$1.5 × more memory than gradient-free methods. When producing the same number of aligned samples, evolutionary methods achieve 3.5×3.5$\times$3.5 × to $\times5$ lower runtime per optimization step compared to gradient-based methods, and only $\times1$ to 1.6×1.6$\times$1.6 × higher running time per step than gradient-free methods. Moreover, evolutionary methods reach higher alignment scores than gradient-free and gradient-based methods in 10 optimization steps. In sum, we contribute:
@@
  • We propose a novel inference-time alignment approach using evolutionary algorithms (EAs). Specifically, we introduce two black-box methods for aligning the outputs of diffusion models: (1) optimizing noise, and (2) optimizing transformations applied to the noise. We elaborate upon the requisite choices for utilizing EAs for inference-time alignment, such as search algorithms, initialization, operators, and computational considerations.
  • We demonstrate the effectiveness of two widely used classes of evolutionary algorithms on the diffusion model alignment task: genetic algorithms and evolutionary strategies. Our evaluation spans a diverse set of reward functions and diffusion models.
2 Related Work
@@
In this section, we review prior work on diffusion model alignment (§ 2.1) and evolutionary algorithms (§ 2.2). An extended treatment of related works can be found in § B.1.
@@ 2.1 Diffusion Model Alignment
Diffusion models use a reverse diffusion process to convert some latent noise distribution into a data distribution, such as images [40, 16]. The reverse diffusion process iteratively denoises the initial latent noise $z_{T}$ over some number of steps $T$ to yield a sample, $z_{0}$. Though diffusion models are capable of modeling complex data distributions, they often fail to produce samples that meet some downstream objective. Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion. These objectives commonly focus on producing images that reflect human aesthetic preferences [17, 51, 52, 53], or other metrics such as compressibility [2].
Alignment methods can generally be grouped into two main categories: fine-tuning–based methods and inference-time methods. Fine-tuning-based alignment methods involve adjusting the diffusion model’s parameters so that generated samples better match alignment objectives [21, 5, 31, 4, 48, 2]. Fine-tuning based methods, although powerful, require retraining and thus all the associated costs.

### v2

# Inference-Time Alignment of Diffusion Models via Evolutionary Algorithms

@@ Abstract
Diffusion models are state-of-the-art generative models, yet their samples often fail to satisfy application objectives such as safety constraints or domain-specific validity. Existing techniques for alignment require gradients, internal model access, or large computational budgets — resulting in high compute demands, or lack of support for certain objectives. In response, we introduce an inference‑time alignment framework based on evolutionary algorithms. We treat diffusion models as black-boxes and search their latent space to maximize alignment objectives. Given equal or less running time, our method achieves 3-35% higher ImageReward scores than gradient-free and gradient-based methods. On the Open Image Preferences dataset, our methods achieve competitive results across four popular alignment objectives. In terms of computational efficiency, we require 55% to 76% less GPU memory and are 72% to 80% faster than gradient-based methods.
footnotetext: * P. Jajal and N.J. Eliopoulos contributed equally to this work.
@@ 1 Introduction
Figure 1: Samples generated by our method on Stable Diffusion-3. Each row shows the progression over optimization steps $i$, with the corresponding metric values displayed in the top-right corner. The noise in each generation is optimized via our evolutionary-based approach, which uses the CoSyNE algorithm [gomez2008accelerated] over 14 optimization steps. Arrows ($\uparrow$$\downarrow$) indicate whether the metric is being maximized or minimized.
@@
Diffusion models [sohl-dickstein_deep_2015, rombach2022high, esser2024scaling] are state-of-the-art generative models for synthesizing high-quality images, video, and audio [chen2023pixart, ho2022video_diffuion, kong2020diffwave, liu2024latent]. However, diffusion model outputs often fail to meet downstream objectives, such as user preferences [xu2023imagereward, hpsv2wu2023human, laionAestheticDataset], safety constraints [miao2024t2vsafetybench], or domain-specific validity [gu2024aligning, wang2024fine]. This problem is referred to as alignment—ensuring generated samples satisfy objectives beyond the original maximum-likelihood objective [liu2024alignmentdiffusionmodelsfundamentals, wallace2024diffusion, black2023training, lee2023aligning, dai2023emu, prabhudesai2023aligning, clark2023directly].
We reiterate the desirable qualities of a diffusion alignment method [uehara2025inference, tang2024inference_dno]: (1) black-box, meaning it applies on a wide range of models because access to gradients or internal model states is not required; (2) support for arbitrary alignment objectives, (3) maintaining sample-efficiency, meaning an objective can be optimized with few samples. Most prior works capture only some of these properties [uehara2025inference]. The works that satisfy all these properties (Best-of-N, Zero-Order) are often used as simple baselines, and we find that these baselines can be competitive with white-box methods. This suggests that alternate, black-box alignment formulations are worth investigating.
In this work, we ask whether black-box evolutionary algorithms (EAs) can be used for diffusion model alignment. EAs should satisfy [eiben2015introduction] the aforementioned criteria, though they have not been applied to the diffusion alignment task. We treat the diffusion model as a black box, and evolve its latent noise vector with EAs to maximize alignment objectives. This approach enables efficient alignment supporting both differentiable and non-differentiable rewards, and is compatible with various diffusion models. Fig.˜1 depicts our method applied to various alignment objectives.
We evaluate our method in terms of its ability to achieve higher rewards than other methods under short-term optimization and equal-compute regimes. Our method is evaluated with two prompt datasets [drawbench2022, data_is_better_together_open_image_preferences_v1_flux_dev_lora_2024], four objective functions [xu2023imagereward, hpsv2wu2023human, clip2021, clark2023directly] (differentiable and non-differentiable), and on Stable Diffusion 1.5 [rombach2022high] as in prior work. Our method outperforms existing gradient-free and gradient-based inference-time alignment approaches on both prompt datasets. We achieve better sample-efficiency, rewards, and memory efficiency over short optimization horizons and under equal-compute regimes. Given similar or less running time, our method achieves 3-35% higher ImageReward scores than gradient-free and gradient-based methods. On the Open Image Preferences dataset, our methods achieve the highest rewards across four popular alignment objectives. In terms of computational efficiency, we require 55% to 76% lower GPU memory and are 72% to 80% faster than gradient-based methods, while consuming up to 1.51.5x more memory than gradient-free methods.
In sum, our contributions are:
@@
  • We propose an inference-time alignment approach using evolutionary algorithms (EAs). Specifically, we introduce two black-box methods for aligning the outputs of diffusion models: (1) directly optimizing latent noise vectors, and (2) optimizing transformations applied to the noise vector. We elaborate on the considerations in using EAs for inference-time alignment, such as search algorithms, initialization, operators, and computational considerations.
  • We evaluate two representative families of EAs (genetic algorithms and evolutionary strategies) for diffusion model alignment. Our evaluation shows generalizability over four popular reward functions, demonstrating EAs are both sample- and computationally-efficient.
2 Background and Related Work
@@
In this section, we review prior work on diffusion model alignment (Sec.˜2.1) and evolutionary algorithms (Sec.˜2.2). Sec.˜B.1 gives an extended treatment of related works.
@@ 2.1 Diffusion Model Alignment
Diffusion models use a reverse diffusion process to convert a latent noise distribution into a data distribution, such as images [sohl-dickstein_deep_2015, ho2022video_diffuion]. The reverse diffusion process (sampling) iteratively denoises the initial latent noise $z_{T}$ over some number of steps $T$ to yield a sample, $z_{0}$. This denoising is guided by conditioning on some variable, usually a text-based prompt. Despite this, they may fail to produce samples that meet some downstream objective. Diffusion model alignment methods adjust diffusion models such that the resulting samples better meet an objective beyond the model’s original maximum likelihood criterion. Examples include aesthetics [kirstain2023pick, hpsv2wu2023human, hpswu2023human, xu2023imagereward], or compressibility [black2023training].